OpenStudy (anonymous):

The sum of a negative number and its square is 72. Find the number. -8 -9 9 8

OpenStudy (anonymous):

x^2 + x = 72 x^2 + x - 72 = 0 (x + 9)(x - 8) = 0 x = -9 and 8 Since the number has to be negative, it is -9.